### v3

@@ 5.4 Acceleration Strategies
We use two heuristics that initially reduce the graph size and set of initial labels and gradually increase them when necessary. The first heuristic consists of separating the set of possible depot leave times (19) into bins {1,…,B} of equal size and sorting these bins in an ascending order with respect to the minimum reduced cost of all labels inside the bin. We then solve the ESPPRC considering only the initial labels in bin $1$. If no column with negative reduced costs has been found, we continue with the next bin until we either find a negative column or solve the final bin $B$. In our algorithm, we set the value $B$ such that each bin $\{1,\dots,B-1\}$ has a size of $5$.
[…]
In our experiments, we set $\Delta=4$ as this returned the best results. Furthermore, we first iterate through all initial label bins and then switch to allowing extensions along all arcs only if no negative column has been found in any of the bins.
@@
[…]
Moreover, Righini and Salani (2008) and Boland, Dethridge, and Dumitrescu (2006) present several other methods to solve the ESPPRC, which rely on relaxing task resources and only adding them if needed. Righini and Salani (2008) develop a branch-and-bound framework, while Boland, Dethridge, and Dumitrescu (2006) propose several algorithms to forbid node repetitions. We note that in the test instances studied, time windows are often so tight that we rarely observed cycles, hence implementing such advanced methods to deal with non-elementary paths does not have a significant impact on the algorithm’s performance for the problem at hand.
@@ 5.5 Branching Strategies
Gélinas et al. (1995) were among the first ones to branch on resource windows, more specifically, time windows. While this strategy is not novel, the process of selecting a task and a time instant to branch on has a significant impact on the branching rule’s efficiency. Let $(\bar{\lambda}^{r})_{r\in\mathcal{R}}$ be an optimal solution to the ARMP and let $\mathcal{R}_{\mathrm{F}}\subseteq\mathcal{R}$ be the set of columns for which $\bar{\lambda}^{r}$ is fractional. For each task $i\in\mathcal{I}$, we define
$\mathcal{B}^{\mathrm{F}}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r\right\},$
$\mathcal{B}^{\mathrm{F}}_{U}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r,\ F_{i}^{r,q}(\omega_{\mathrm{max}})\ \text{unique}\right\}$
as the set of all (unique) worst-case finish times of task $i$ within team routes with fractional values, respectively. We then select all tasks for which the cardinality of the latter set is maximal and denote it by $\mathcal{I}^{\mathrm{F}}$, i.e.,
@@
We then select the task $i^{*}$ for which the standard deviation of the worst-case finish times of task $i^{*}$ in all routes in $\mathcal{B}^{\mathrm{F}}(i)$ is maximal, that is
[…]
@@
$F_{i^{*}}^{r,q}(\omega_{\mathrm{max}})\leq\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r\quad\text{or}\quad F_{i^{*}}^{r,q}(\omega_{\mathrm{max}})>\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r,$
@@
$\max\left\{\sigma\left(\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{B}^{\mathrm{F}}(i)\right\}\right):\ i\in\mathcal{I}^{\mathrm{F}}\right\}=0$
@@
Desrochers, Desrosiers, and Solomon (1992) introduced branching on tour counts, which has proven to be efficient for vehicle routing problems. Let $\tau^{*}$ be the time instant for which $l_{\tau^{*}}=\underset{(r,q)\in\mathcal{R}}{\sum}g_{\tau^{*}}^{r}\lambda^{r}_{q}$ is closest to 0.5, where $g_{\tau}^{r}=1$ if $\mathrm{tl}^{r,q}\leq g_{\tau}^{r}\leq\mathrm{tr}^{r,q}$ and $g_{\tau}^{r}=0$ else. We then branch on the current solution by imposing
@@
The pricing problem has to be adjusted in the sense that time-dynamic arc weights have to consider the dual cost of the additional inequalities. If a label $L$ is extended along an arc $(v,v^{\prime})$ and $F_{v^{\prime}}^{L}(\omega_{\mathrm{max}})<\tau^{*}\leq F_{v^{\prime}}^{L}(\omega_{\mathrm{max}})$ holds, the cost of arc $(v,v^{\prime})$ must be adjusted by the dual cost of (25). Similar to branching on task finish times, this branching strategy can potentially produce a trivial branch, making a fallback branching option necessary.
@@ 5.6 Cutting Planes
Whenever a fractional optimal solution is found at a node in the branching tree, we model the problem of finding a most violated rank-1 Chvátal-Gomory cut (CGC) as a mixed-integer problem and solve it using a generic solver with a very short time limit. If optimality could not be proven within the time limit, the best solution found so far is returned. If no solution is found, the cutting plane step is skipped and the algorithm proceeds with a branching step. This approach was first proposed by Fischetti and Lodi (2007) and applied to a vehicle routing problem with time windows by Petersen, Pisinger, and Spoorendonk (2008). In the following, we base our considerations on the ARMP formulation. Let $(u_{i})_{i\in\mathcal{I}}\in[0,1)$ and $(u_{k,t})_{k\in\mathcal{K},t\in\mathcal{T}}$ be the coefficients of the most violated cut. We then add the constraint
$\underset{(r,q)\in\mathcal{R}}{\sum}\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})u_{k,t}\right\rfloor\lambda^{r}_{q}\leq\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}N_{k}\cdot\left(\underset{\tau\in\mathcal{T}}{\sum}u_{k,t}\right)\right\rfloor$
@@
[…]
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}\right\rfloor\cdot\psi_{g}$
$T_{v^{\prime},g}$ $=T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}-\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}\right\rfloor$
where $\psi_{g}\geq 0$ is the dual variable associated with CGC $g\in\mathcal{G}$. Dominance rule 5.1 is then slightly adjusted and extended:
Definition 5.4 (Dominance Rule)
@@
[…]
@@
Inequalities (26) and (27) ensure that for any feasible extension $(v,v^{\prime})$ of $L^{2}$, the reduced costs still remain dominated by $L^{1}$. For a proof of correctness of this approach, we refer to Section 4 of Petersen, Pisinger, and Spoorendonk (2008).
@@ 6 Experimental Study
Section 6.1 elaborates on the generation of test instances. Section 6.2 summarizes the algorithm’s performance for different configurations of the aforementioned components. In Section 6.3, stochastic and deterministic optimal policies are compared with respect to their practical feasibility. The algorithm is implemented using Python 3.11 and Gurobi 12.0.0. Furthermore, all computational studies were performed on a single machine equipped with an Intel® Xeon® W-1390p 11th gen 8-core 3.5GHz processor, 32GB of RAM and running Windows 10. All instances and corresponding solutions can be found under https://github.com/andreashagntum/StochasticTeamFormationRoutingAirport.

### v4

@@ 5.4 Acceleration Strategies
We use two heuristics that initially reduce the graph size and the set of initial labels and gradually increase them when necessary. The first heuristic consists of separating the set of possible depot leave times into containers {1,…,C} of equal size and sorting these containers in an ascending order with respect to the minimum reduced cost of all labels inside the container. We then solve the ESPPRC considering only the initial labels in container $1$. If no column with negative reduced costs has been found, we continue with the next container until we either find a negative column or solve the final container $C$. In our algorithm, we set the value $C$ such that each container $\{1,\dots,C-1\}$ has a size of $5$.
The second heuristic is similar to the one proposed by desaulniers2006tabu. For each task node $i$, we sort the set of outgoing arcs $(i,j)\in\mathcal{A}^{q}$ with $j\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}$ based on their task duals $\mu_{j}$. Initially, we only allow extensions along $(i,j)$ if $\mu_{j}$ is among the $\Delta$ largest dual values of task nodes adjacent to $i$. If we do not find a negative column, we solve the pricing problem while allowing extensions along all arcs.
In our experiments, we set $\Delta=4$ as this returned the best results. Furthermore, we first iterate through all initial label containers and then switch to allowing extensions along all arcs only if no negative column has been found in any of the containers.
@@
Furthermore, righini2008new propose directional labeling, during which labels are also extended backwards and eventually joined with forward labels, which often greatly reduces the number of explored labels. In order to apply this technique to the problem at hand, when backwards-extending a label $L\coloneqq\left(\mathrm{tl}^{L},v,P,T_{v}^{cost},(T_{v,i}^{perf})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},F_{v}^{L},\right)$ along $(v^{\prime},v)$, it needs to be possible to derive the finish time distribution $F_{v^{\prime}}^{L^{\prime}}$ based on $F_{v}^{L}$ and the distribution of travel times $\left(\mathbbm{P}(t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}_{v^{\prime},v}\leq\tau)\right)_{\tau\in B_{i,j}}$. However, when distributions are left-truncated as in the case of hard time windows, this is generally not the case (see G). Alternatively, one can extend the node set of the graph such that distributional properties that are relevant for the stochasticity considered are embedded in the nodes. For instance, GAUVIN2014141 consider stochastic demand and minimize routing costs that consist of deterministic travel costs and expected costs of failure, which in turn depend on the cumulative demand distribution at each customer. The authors assume independent, Poisson-distributed demands. Hence, the cumulative demand distribution is again Poisson-distributed. This allows them to define a graph that contains nodes for each pair of customer and distributional parameter, i.e., the expected cumulative demand at the current customer. We note that this approach heavily relies on the possibility of propagating distributional properties, in this case, a Poisson distribution, along vehicle routes. Once distributions are truncated, as is the case for hard time windows, this is generally no longer possible.
Moreover, righini2008new and BOLAND200658 present several other methods to solve the ESPPRC, which rely on relaxing task resources and only adding them if needed. righini2008new develop a branch-and-bound framework, while BOLAND200658 propose several algorithms to forbid node repetitions. We note that in the test instances studied, time windows are often so tight that we rarely observed cycles, hence implementing such advanced methods to deal with non-elementary paths does not have a significant impact on the algorithm’s performance for the problem at hand.
@@ 5.5 Branching Strategies
gelinas1995new were among the first ones to branch on resource windows, more specifically, time windows. While this strategy is not novel, the process of selecting a task and a time instant to branch on has a significant impact on the branching rule’s efficiency. Let $(\bar{\lambda}^{r})_{r\in\mathcal{R}}$ be an optimal solution to the ARMP and let $\mathcal{R}_{\mathrm{F}}\subseteq\mathcal{R}$ be the set of columns for which $\bar{\lambda}^{r}$ is fractional. For each task $i\in\mathcal{I}$, we define
$\mathcal{B}^{\mathrm{F}}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r\right\},$
$\mathcal{B}^{\mathrm{F}}_{U}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r,\ F_{i}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})\ \text{unique}\right\}$
as the set of all (unique) $\omega_{\gamma}$-scenario finish times of task $i$ within team routes with fractional values, respectively. We then select all tasks for which the cardinality of the latter set is maximal and denote it by $\mathcal{I}^{\mathrm{F}}$, i.e.,
@@
We then select the task $i^{*}$ for which the standard deviation of the $\omega_{\gamma}$-scenario finish times of task $i^{*}$ in all routes in $\mathcal{B}^{\mathrm{F}}(i)$ is maximal, that is
$i^{*}=\operatorname*{arg\,max}\left\{\sigma\left(\left\{F_{i}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}):\ (r,q)\in\mathcal{B}^{\mathrm{F}}(i)\right\}\right):\ i\in\mathcal{I}^{\mathrm{F}}\right\}$
@@
$F_{i^{*}}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})\leq\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r\quad\text{or}\quad F_{i^{*}}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})>\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r,$
@@
$\max\left\{\sigma\left(\left\{F_{i}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}}):\ (r,q)\in\mathcal{B}^{\mathrm{F}}(i)\right\}\right):\ i\in\mathcal{I}^{\mathrm{F}}\right\}=0$
@@
desrochers_branch_on_vehicle_count introduced branching on tour counts, which has proven to be efficient for vehicle routing problems. Let $\tau^{*}$ be the time instant for which $l_{\tau^{*}}=\underset{(r,q)\in\mathcal{R}}{\sum}g_{\tau^{*}}^{r}\lambda^{r}_{q}$ is closest to 0.5, where $g_{\tau}^{r}=1$ if $\mathrm{tl}^{r,q}\leq g_{\tau}^{r}\leq\mathrm{tr}^{r,q}$ and $g_{\tau}^{r}=0$ else. We then branch on the current solution by imposing
@@
The pricing problem has to be adjusted in the sense that time-dynamic arc weights have to consider the dual cost of the additional inequalities. If a label $L$ is extended along an arc $(v,v^{\prime})$ and $F_{v^{\prime}}^{L}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})<\tau^{*}\leq F_{v^{\prime}}^{L}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})$ holds, the cost of arc $(v,v^{\prime})$ must be adjusted by the dual cost of (27). Similar to branching on task finish times, this branching strategy can potentially produce a trivial branch, making a fallback branching option necessary.
@@ 5.6 Cutting Planes
Whenever a fractional optimal solution is found at a node in the branching tree, we model the problem of finding a most violated rank-1 Chvátal-Gomory cut (CGC) as a mixed-integer problem and solve it using a generic solver with a very short time limit. If optimality could not be proven within the time limit, the best solution found so far is returned. If no solution is found, the cutting plane step is skipped and the algorithm proceeds with a branching step. This approach was first proposed by fischetti2007optimizing and applied to a vehicle routing problem with time windows by petersen2008chvatal. In the following, we base our considerations on the ARMP formulation. Let $(u_{i})_{i\in\mathcal{I}}\in[0,1)$ and $(u_{k,t})_{k\in\mathcal{K},t\in\mathcal{T}}$ be the coefficients of the most violated cut. We then add the constraint
$\underset{(r,q)\in\mathcal{R}}{\sum}\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}b_{k,\tau}^{r,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})u_{k,t}\right\rfloor\lambda^{r}_{q}\leq\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}N_{k}\cdot\left(\underset{\tau\in\mathcal{T}}{\sum}u_{k,t}\right)\right\rfloor$
@@
Let $\mathcal{G}$ be the set of indices of CGCs and $\mathcal{C}^{\mathcal{G}}\coloneqq\{\left((u_{i}^{g})_{i\in\mathcal{I}},(u_{k,t}^{g})_{k\in\mathcal{K},t\in\mathcal{T}}\right):\ g\in\mathcal{G}\}$ be the cut coefficients for all CGCs present at the current node. During the pricing step, we introduce an additional resource $T_{v,g}$ for each CGC $g\in\mathcal{G}$. When extending a label $L\coloneqq\left(\mathrm{tl}^{L},v,P,T_{v}^{\mathrm{cost}},(T_{v,i}^{\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},(T_{v,g})_{g\in\mathcal{G}},\mathcal{F}_{F_{v}^{L}}\right)$ along an arc $(v,v^{\prime})$, we update the resources $(T_{v,g})_{g\in\mathcal{G}}$ and $T_{v^{\prime}}^{\mathrm{cost}}$ using the following resource extension function:
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\overset{F_{v^{\prime}}^{L^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}b_{k,\tau}^{P,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})u_{k,t}^{g}\right\rfloor\cdot\psi_{g}$
$T_{v^{\prime},g}$ $=T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\overset{F_{v^{\prime}}^{L^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}b_{k,\tau}^{P,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})u_{k,t}^{g}-\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\overset{F_{v^{\prime}}^{L^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\omega_{\gamma})}}{\sum}}b_{k,\tau}^{P,q}(\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma}})u_{k,t}^{g}\right\rfloor$
where $\psi_{g}\geq 0$ is the dual variable associated with CGC $g\in\mathcal{G}$. Dominance rule 1 is then slightly adjusted and extended:
Definition 3 (Dominance Rule).
@@
$L^{k}\coloneqq\left(\mathrm{tl}^{k},v,P^{k},T_{v}^{k,\mathrm{cost}},(T_{v,i}^{k,\mathrm{perf}})_{i\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}_{q}},(T_{v,g}^{k})_{g\in\mathcal{G}},F_{v}^{k}\right)$
@@
Inequalities (28) and (29) ensure that for any feasible extension $(v,v^{\prime})$ of $L^{2}$, the reduced costs still remain dominated by $L^{1}$. For a proof of correctness of this approach, we refer to Section 4 of petersen2008chvatal.
@@ 6 Experimental Study
Section 6.1 elaborates on the generation of test instances. Section 6.2 summarizes the algorithm’s performance for different configurations of the aforementioned components, as well as a performance comparison to the approach proposed by yuan2015branch. In Section 6.3, stochastic and deterministic optimal policies are compared with respect to their practical feasibility. The proposed algorithmm as well as the benchmark approach by yuan2015branch, were implemented using Python 3.12 and Gurobi 12.0.2. Furthermore, all computational studies were performed on a single machine equipped with an Intel® Core® i7-8700K 11th 12-core 3.7GHz processor, 32GB of RAM and running Windows 11. All instances and corresponding solutions can be found under https://github.com/andreashagntum/StochasticTeamFormationRoutingAirport.